OpenStudy (rainbow_rocks03):

The regular price of an item is $65. The item is on sale with a discount of $8. What is the discount rate? A. 0.123% B. 1.23% C. 8% D. 12.3%

OpenStudy (bloomlocke367):

okay.. what you can do is write a proportion..

OpenStudy (bloomlocke367):

\[\frac{ 8 }{ 65 }=\frac{ ? }{ 100 }\] do you understand why I set it up that way? cross multiply and divide to find the '?'

OpenStudy (rainbow_rocks03):

8*?=?

OpenStudy (bloomlocke367):

no.. CROSS multiply and divide XD 8*100= what? then divide that by 65

OpenStudy (rainbow_rocks03):

um i did that and it was a long crazy number.

OpenStudy (bloomlocke367):

800/65?

OpenStudy (rainbow_rocks03):

yah

OpenStudy (bloomlocke367):

you should've gotten 12.307.... what does that round to?

OpenStudy (rainbow_rocks03):

12.3?

OpenStudy (bloomlocke367):

yes.. :) and that's already the percent because it was over 100

OpenStudy (rainbow_rocks03):

so 12.3 is my answer.

OpenStudy (bloomlocke367):

yes. 12.3%